OpenStudy (anonymous):

T(n) is the total number of lights in a triangular lighting rig of n rows, as shown. Which polynomial function represents T(n + 1)? T(n)=1/2n^2+1/2n

OpenStudy (experimentx):

put n+1 instead of n in T(n)=1/2n^2+1/2n

OpenStudy (anonymous):

T(n+1)=1/2n^2+1/2(n+1)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

??

OpenStudy (turingtest):

T(n+1)=1/2n^2+1/2(n+1) ^ you forgot to put n+1 here as well

OpenStudy (anonymous):

so wat wud b tha awnser

OpenStudy (turingtest):

the answer would be what you wrote above, but with n+1 in place of n

OpenStudy (turingtest):

...all the n's in 1/2n^2+1/2n get replaced by n+1 you did it right for the second n, but you forgot to do it for the first

OpenStudy (anonymous):

a. T(n+1)=1/2n^2+3/2n+1 b. T(n+1)=1/2n^2+1/2n+1 c. T(n+1)=1/2n^2+n+1 d. T(n+1)=1/2n^2+1 ??

OpenStudy (turingtest):

I'm not just going to give you the answer that is against the code of conduct here

OpenStudy (turingtest):

however I'm not so sure the answer is on the list lol looks like you have to do some simplification first